Ron had a square and changed it's side by adding one unit to one pair of opposite sides and shortening the other pair by the same lenght. As a result, he obtained a rectangle with an area of 80 units. With what square did he start?
Thank you very much , if you are able to help.

Answer by stanbon(75887)

Assume the side of the original square figure was x.
The new figure would have dimensions (x-1) and (x+1)
So Area= (x+1)(x-1)=80
x^2-1=80
x^2=81
x=9
The original square has a side of 9 units.
Cheers,
Stan H.